OpenStudy (anonymous):

A rubber ball dropped on a hard surface takes a sequence of bounces, each one 3/5 as high as the preceding one. If this ball is dropped from a height of 10 feet, what is the total vertical distance it has traveled after it hits the surface the 5th time?

Parth (parthkohli):

Add this 10 + 10(3/5) + 10(3/5)^2 + 10(3/5)^3 + 10(3/5)^4 + 10(3/5)^5

OpenStudy (anonymous):

@ParthKohli dont got my calc on me :/

Parth (parthkohli):

Use your brain then :D

Parth (parthkohli):

Or use wolfram..

OpenStudy (campbell_st):

there are 2 series here.... the down series where a = 10 and r = 3/5 and n = 5 the up series where a = 6 and r = 3/5 and n = 4 as after bouncing the ball only goes up 4 times this is a geometric series where r < 1 so use \[s _{n}= \frac {a(1 - r^n)}{1 - r}\] \[s _{5}= \frac{10(1 - (\frac{3}{5})^5)}{1 - \frac{3}{5}} + \frac{6(1 - (\frac{3}{5})^4}{1 - \frac{3}{5}}\] just evaluate it